A firm has analyzed its operating conditions, prices, and cost have developed the following functions: R = 400 – 4q2 (per thousand naira) and Cost = q2 + 10q + 30 (per thousand naira) where q is the number of units produced and sold. If the firm wishes to maximize profit:

(a) What quantity should be sold and at what price?

(b) What will be the amount of profit?

2. Find the point of maximum value of the revenue function given as P = 400 – 4q. Hence, find the revenue.



1
Expert's answer
2021-08-25T17:07:28-0400

1.

(a) "R= 400-4Q^2"

"MR=8Q"

"C=Q^2+10Q+30"

"MC=2Q+10"

At maximum profit:

"MR=MC"

"8Q=2Q+10"

"Q= 1.66667"

Q is approximately =2

(b)

"R=400-4(4)=384"

"C=4+20+30=54"

Profit"= 384-54=330"


2.

"P=400-4Q"

Max value of revenue is 4

Revenue "= 400-4(4)=384"